# LEADRE: Multi-Faceted Knowledge Enhanced LLM Empowered Display Advertisement Recommender System

### v1

@@ Abstract.
Display advertising provides significant value to advertisers, publishers, and users. Traditional display advertising systems utilize a multi-stage architecture consisting of retrieval, coarse ranking, and final ranking. However, conventional retrieval methods rely on ID-based learning to rank mechanisms and fail to adequately utilize the content information of ads, which hampers their ability to provide diverse recommendation lists.
To address this limitation, we propose leveraging the extensive world knowledge of LLMs. However, three key challenges arise when attempting to maximize the effectiveness of LLMs: ”How to capture user interests”, ”How to bridge the knowledge gap between LLMs and advertising system”, and ”How to efficiently deploy LLMs”. To overcome these challenges, we introduce a novel LLM-based framework called LLM Empowered Display ADvertisement REcommender system (LEADRE). LEADRE consists of three core modules: (1) The Intent-Aware Prompt Engineering introduces multi-faceted knowledge and designs intent-aware ¡Prompt, Response¿ pairs that fine-tune LLMs to generate ads tailored to users’ personal interests. (2) The Advertising-Specific Knowledge Alignment incorporates auxiliary fine-tuning tasks and Direct Preference Optimization (DPO) to align LLMs with ad semantic and business value. (3) The Efficient System Deployment deploys LEADRE in an online environment by integrating both latency-tolerant and latency-sensitive service. Extensive offline experiments demonstrate the effectiveness of LEADRE and validate the contributions of individual modules. Online A/B test shows that LEADRE leads to a 1.57% and 1.17% GMV lift for serviced users on WeChat Channels and Moments separately. LEADRE has been deployed on both platforms, serving tens of billions of requests each day.
Display Advertising, LLM-based Recommendation, Generative Retrieval
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Display advertisingccs: Information systems Personalization
@@ 1. Introduction
Online display advertising provides significant value to advertisers, publishers, and users by facilitating targeted content delivery and meeting users’ personal interests (Song et al., 2024; Singh et al., 2023). Traditionally, display advertising systems utilize a multi-stage architecture, including retrieval, coarse ranking, ranking, re-ranking, etc. The retrieval stage, serving as the entry of the process, is crucial for identifying user interests and overcoming the ”information cocoon” by presenting a diverse set of ad options(Li et al., 2021; Chen et al., 2022).
@@
Traditional retrieval solutions typically employ ID-based learning-to-rank mechanisms to learn collaborative semantics for efficient ad filtering. However, these methods fail to adequately utilize the content information of ads, bearing limitations in generating diverse recommendations—especially when dealing with sparse user behaviors and long-tail ads (Wei et al., 2021; Dong et al., 2020).
@@
Recently, large language models (LLMs) have demonstrated exceptional capabilities in understanding, generalization, and reasoning by leveraging vast amounts of general knowledge (Nam et al., 2024; Zhang et al., 2024a; Li and Zhang, 2024). To address the shortcomings of conventional retrieval approaches, researchers have begun incorporating LLMs into recommender systems (Rajput et al., 2024; Wei et al., 2024). However, these methods often struggle to find applications in industrial display advertising systems, which introduces several critical challenges:
@@
(1) How to capture user interests in scenarios with implicit user intents and sparse user behaviors: While LLMs have demonstrated strong capabilities in understanding user intent, display advertising often operates in an environment with implicit user intents, where explicit queries are absent. Additionally, user behavior in the ad domain tends to be sparse, making it crucial to effectively utilize the limited available data and supplement it appropriately to capture commercial intent. (2) How to bridge the knowledge gap within LLMs and advertising system: Although LLMs excel at generating natural language responses based on general knowledge, display advertising systems require the generation of specific ads from a predefined set of ads. Bridging the gap between the general capabilities of LLMs and the specific needs of the advertising domain is essential. Moreover, the generated ads must align with business objectives, necessitating that LLMs are tuned to meet the specific goals of the advertising system. (3) How to efficiently deploy LLMs in industrial display advertising system which serves billions of users requests per day: Deploying LLMs within an online advertising system introduces significant computational overhead, which can conflict with the need to maintain cost efficiency. Therefore, it is crucial to deploy LLM-based methods in a way that balances computational costs with advertising performance, ensuring scalability and efficiency within the system.
@@
To address these challenges, we propose a novel multi-faceted knowledge enhanced LLM Empowered display ADvertisement REcommender system (LEADRE), termed LEADRE. The framework consists of three key modules:
@@
(1) Intent-Aware Prompt Engineering: This module designs intent-aware ¡Prompt, Response¿ pairs that fine-tune LLMs to generate ads tailored to users’ personal interests. To overcome the scarcity of user behaviors in the ad domain, this module introduces user behaviors from content domains such as micro-videos, and news. Additionally, it models user commercial intent by combining long-term interests (derived from user profiles and historical behaviors) and short-term interests (derived from recent behaviors) within the prompt. (2) Advertising-Specific Knowledge Alignment: This module incorporates auxiliary fine-tuning tasks and Direct Preference Optimization (DPO). The auxiliary fine-tuning tasks are specifically designed to bridge the semantic gap between language modality and advertising system. Additionally, to balance user intent and business value, we apply DPO, which encourages the model to generate ads with higher business value. (3) Efficient System Deployment: This module integrates both latency tolerant service and latency sensitive service into the LLM deployment architecture. To further improve computational efficiency, we optimize the deployment process using TensorRT LLM Acceleration.
@@
We implement LEADRE using Hunyuan with 1B parameters and conduct extensive offline and online experiments within Tencent’s display advertising system. The offline results validate the contributions of the individual modules, demonstrating their effectiveness in the overall framework. In the online A/B test, LEADRE achieved a 1.57% increase in Gross Merchandise Value (GMV) on Tencent WeChat Channels for serviced users, along with a 1.17% increase in GMV on Tencent WeChat Moments for serviced users. Currently, LEADRE has been successfully deployed on both platforms, serving billions of users and processing tens of billions of user requests each day. The retrieved ads are further incorporated into the ranking phase as new features. Specifically, on the user side, the retrieved ads are used to extended user interests, while on the item side, the match scores between the retrieved ads and the target ad are introduced as new item-level features. These new features contribute to an extra 1.43% improvement in GMV on Tencent WeChat Channels.
@@
(1) To the best of our knowledge, this is the first work to deploy LLMs in an online display advertising system. We propose a novel LLM-based generative retrieval framework, LEADRE, for display advertising, and deploy it using a combination of latency tolerant service system and latency sensitive service system to optimize performance and resource utilization.
(2) LEADRE integrates an Intent-Aware Prompt Engineering and an Advertising-Specific Knowledge Alignment to ensure the generated ads are accurate, diverse, aligned with high business value, and meet user interests.
(3) Extensive experiments, both offline and online, demonstrate the effectiveness of the proposed method. We observe significant improvements in offline metrics such as HitRatio, as well as in online metrics like GMV.
@@ 2.1. Large Language Models-based Recommender
Large Language Models (LLMs) have demonstrated remarkable capabilities in both understanding and generation tasks(Yao et al., 2024; Huang et al., 2022), and they have been widely applied to a variety of domains, including document summarization(Koh et al., 2022; Zhang et al., 2024b), conversational agents(He et al., 2023; Wang et al., 2023), code completion(Guo et al., 2023; Zhang et al., 2024c), and others(Zhu et al., 2023; Zhai, 2024). Recently, researchers have begun exploring the application of LLMs in recommender systems(Wu et al., 2024). These efforts can be broadly categorized into two roles for LLMs: feature encoder and ranker. As feature encoders, LLMs are leveraged primarily for their ability to understand and represent textual information. In these works, user or item features are often assembled into textual descriptions via pre-designed templates, which are then encoded with LLMs to generate feature embeddings (Ren et al., 2024; Wei et al., 2024; Bao et al., 2024). These embeddings can subsequently be utilized in downstream tasks, such as ranking items in a recommender system. On the other hand, as rankers, LLMs are employed for their ability to generate. In these works, user behavior sequences are transformed into text sequences, and LLMs are tasked with generating the next likely item (or user action) based on the input text (Zhang et al., 2023; Harte et al., 2023). To this end, LLMs typically require fine-tuning on recommendation-specific tasks to acquire the domain knowledge and capture user preference in recommender systems (Bao et al., 2023; Harte et al., 2023; Lin et al., 2024).
However, these works fail to find applications in industrial display advertising systems due to implicit user intents, high computational costs, etc. In this work, we present the first application of LLM-based generative retrieval in an industrial display advertising system. We employ intent-aware prompt engineering and ad-specific knowledge alignment to enable LLMs to generate ads that are diverse, aligned with business objectives, and tailored to user interests. Additionally, we address computational cost challenges by deploying LLMs through a hybrid architecture that integrates both latency-tolerant and latency-sensitive services.
@@ 2.2. Retrieval in Advertising Systems
Display advertising systems typically follow a multi-stage architecture composed of multiple stages, including retrieval, corse-ranking, ranking, re-ranking, and others(Cui et al., 2020; Gharibshah et al., 2020). The retrieval stage, positioned at the top of the funnel, plays a critical role in identifying user interests and generating a pool of candidate ads for further processing (Xie et al., 2022; Ji et al., 2024). Most advertising systems adopt ID-based learning to ranking mechanisms to recall all ads that users may find appealing. These methods often rely on a two-tower model, where one tower encodes user features and the other encodes item (ad) features(Cen et al., 2020; Huang et al., 2020).
Despite their efficiency, ID-based retrieval methods can suffer from issues such as the ”information cocoon” effect, limited diversity in the retrieved list, and the inability to effectively leverage the ads’ content information(Hurley and Zhang, 2011; Wu et al., 2022). In this work, we propose incorporating LLMs to address these issues. LLMs, with their powerful generative and contextual understanding capabilities, can enhance the diversity of the retrieved list by expanding the range of ads presented to users, ultimately improving the overall effectiveness of the advertising system.
@@ 3.1. Problem Definition
Let $\mathcal{U}$ and $\mathcal{A}$ denote the sets of users and advertisements(ads), respectively, within the target advertising system. Due to the limited user interactions with the ads, we introduce a content domain to enrich user behavior data. Let $\mathcal{C}$ denote the set of content items in the system. The sequence of user behaviors is represented as $\mathcal{S}_{u}=[i^{1}_{u},i^{2}_{u},\cdots,i^{L}_{u}]$ in chronological order, where $u\in\mathcal{U}$, $i_{u}\in\mathcal{A}\cup\mathcal{C}$, and $L=|\mathcal{S}_{u}|$ is the length of the user’s behavior sequence. Given this user behavior sequence, the goal of generative retrieval is to generate the next relevant ad, $a^{L+1}_{u}\in\mathcal{A}$.
@@ 3.2. Ads Indexing
Traditional advertising systems typically index ads incrementally and learn dense embeddings for each ad (Wu et al., 2020; Chang et al., 2021). However, the index built in this way usually lacks semantic information, creating a significant gap between the advertising system and human understanding about ads. To bridge this gap, we adopt the strategy of learning Semantic IDs (S-IDs) based on ad features (illustrated in Figure 1), inspired by previous works on item indexing (Rajput et al., 2024; Zheng et al., 2024; Geng et al., 2022).
Specifically, we first define a feature mapping rule in the form of a template string to generate text description of each ad, as follows:
@@
After obtaining the textual description $t_{a}$, we employ a pre-trained language model (e.g. Hunyuan (Sun et al., 2024), E5(Wang et al., 2022)) to encode $t_{a}$, bridging the gap between the advertising system and human understandable text. This step can be formulated as $\mathbf{x}_{a}=f_{PLM}(t_{a};\theta_{PLM})$, where $f_{PLM}$ represents the pre-trained language model parameterized by $\theta_{PLM}$. The resulting $\mathbf{x}_{a}\in\mathbb{R}^{d_{PLM}}$ is the textual embedding of ad $a$, and $d_{PLM}$ denotes the output dimension of the pre-trained language model.
After that, we employ a Residual-Quantized Variational AutoEncoder (RQ-VAE) (Lee et al., 2022) to generate concise semantic IDs for the ads. The RQ-VAE takes the ad’s embedding $\mathbf{x}_{a}$ as input and uses an encoder-quantization-decoder mechanism to generate a list of discrete semantic tokens. For details of RQ-VAE training please refer to Appendix A.1

### v3

@@ Abstract.
Display advertising plays a crucial role in benefiting advertisers, publishers, and users. Traditional display advertising systems employ a multi-stage architecture comprising retrieval, coarse ranking, ranking, and re-ranking. However, conventional retrieval methods primarily rely on ID-based learning-to-rank mechanisms, often underutilizing the content information of ads, like ads’ title, and description. This limitation reduces the ability to generate diverse and relevant recommendation lists.
To address this challenge, we propose leveraging the extensive world knowledge of large language models (LLMs). However, effectively integrating LLMs into advertising systems presents three key challenges: (i) How to accurately capture user interests, (ii) How to bridge the knowledge gap between LLMs and advertising systems, and (iii) How to efficiently deploy LLMs at scale. To overcome these challenges, we introduce LEADRE—the LLM Empowered Display ADvertisement REcommender system. LEADRE consists of three core components. The Intent-Aware Prompt Engineering module introduces multi-faceted knowledge and constructs intent-aware ¡Prompt, Response¿ pairs, fine-tuning LLMs to generate ads tailored to users’ personal interests. The Advertising-Specific Knowledge Alignment module incorporates auxiliary fine-tuning tasks and Direct Preference Optimization (DPO) to align LLMs with advertising semantics and business objectives. The Latency-Aware Model Deployment module integrates a hybrid service framework that balances latency-tolerant and latency-sensitive service, ensuring seamless online deployment.
Extensive offline experiments validate the effectiveness of LEADRE, demonstrating significant improvements across multiple evaluation metrics. Furthermore, online A/B tests reveal a 1.57% and 1.17% increase in Gross Merchandise Value (GMV) for serviced users on WeChat Channels and Moments, respectively. LEADRE has been successfully deployed on both platforms, handling tens of billions of requests daily.
PVLDB Reference Format:
PVLDB, 18(12): XXX-XXX, 2025.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 18, No. 12 ISSN 2150-8097.
doi:XX.XX/XXX.XX
@@ 1. Introduction
Online display advertising plays a crucial role in facilitating targeted content delivery and meeting users’ personal interest, benefiting advertisers, publishers, and users (Song et al., 2024; Singh et al., 2023). Traditional display advertising systems employ a multi-stage architecture, including retrieval, coarse ranking, ranking, and re-ranking. The retrieval stage, which initiates the process, is critical for identifying user interests and mitigating the ”information cocoon” effect by providing a diverse set of ad options (Li et al., 2021; Chen et al., 2022).
@@
Conventional retrieval methods primarily rely on ID-based learning-to-rank mechanisms to learn collaborative semantics for efficient ad filtering. However, these methods often underutilize ad content information, such as the title and description of the ads, which limits in generating diverse recommendations, particularly in scenarios with sparse user behaviors and long-tail ads (Wei et al., 2021; Dong et al., 2020).
@@
Recently, large language models (LLMs) have demonstrated remarkable capabilities in understanding, generalization, and reasoning by leveraging vast amounts of general knowledge (Nam et al., 2024; Zhang et al., 2024a; Li and Zhang, 2024). Researchers have explored incorporating LLMs into recommender systems (Rajput et al., 2024; Wei et al., 2024) to enhance retrieval performance. However, applying LLMs in industrial-scale display advertising presents several critical challenges:
@@
(1) Capturing user interests in scenarios with implicit intents and sparse behaviors: While LLMs exhibit strong capabilities in understanding user intent, display advertising often lacks explicit queries, making it difficult to infer user intent. Additionally, user interactions in the ad domain are typically sparse, necessitating effective utilization of limited data and supplementary behaviors to enhance commercial intent modeling. (2) Bridging the knowledge gap between LLMs and advertising systems: Although LLMs excel in generating natural language responses based on general knowledge, display advertising requires the generation of relevant ads from a predefined inventory. It is essential to bridge the gap between the general knowledge of LLMs and the specific requirements of advertising systems. Furthermore, the generated ads must align with business objectives, requiring tailored fine-tuning strategies to optimize LLMs for advertising-specific goals. (3) Efficiently deploying LLMs in large-scale advertising systems serving billions of requests per day: Integrating LLMs into an online advertising system imposes substantial computational demands, potentially conflicting with cost efficiency requirements. Thus, optimizing LLM deployment is essential to balance computational costs while maintaining scalability and real-time responsiveness.
@@
To address these challenges, we propose LEADRE—a multi-faceted knowledge enhanced LLM Empowered display ADvertisement REcommender system. LEADRE consists of three core components:
@@
(1) Intent-Aware Prompt Engineering: This module designs intent-aware ¡Prompt, Response¿ pairs that fine-tune LLMs to generate ads tailored to users’ interests. To mitigate data sparsity, user behaviors from related content domains, such as micro-videos and news, are integrated. Furthermore, commercial intent is modeled by incorporating both long-term interests (from user profiles and historical behaviors) and short-term interests (from recent interactions) into the prompt. (2) Advertising-Specific Knowledge Alignment: To bridge the semantic gap between natural language and advertising data, this module introduces auxiliary fine-tuning tasks. Additionally, Direct Preference Optimization (DPO) is applied to balance user intent with business objectives, ensuring that the generated ads have high commercial value. (3) Latency-Aware Model Deployment: LEADRE is deployed using a hybrid architecture that integrates both latency-tolerant and latency-sensitive services. To further enhance computational efficiency, we optimize deployment using TensorRT LLM Acceleration.
@@
We implement LEADRE using Hunyuan with 1B parameters and evaluate its performance through extensive offline and online experiments in Tencent’s display advertising system. The offline results validate the contributions of the individual modules, demonstrating their effectiveness. In online A/B tests, LEADRE achieved a 1.57% increase in Gross Merchandise Value (GMV) on Tencent WeChat Channels and a 1.17% increase on Tencent WeChat Moments. Currently, LEADRE is deployed on both platforms, serving billions of users and processing tens of billions of requests per day.
The retrieved ads are further incorporated into the ranking phase as additional features. Specifically, on the user side, retrieved ads extend user interest representations, while on the item side, match scores between retrieved ads and target ads serve as new item-level features. These enhancements contribute to an additional 1.43% improvement in GMV on Tencent WeChat Channels.
@@
  • To the best of our knowledge, this is the first study to deploy LLMs in an online display advertising system. We introduce LEADRE, a novel LLM-based generative retrieval framework, and deploy it through a hybrid architecture that integrates both latency-tolerant and latency-sensitive service.
  • LEADRE integrates Intent-Aware Prompt Engineering and Advertising-Specific Knowledge Alignment to ensure the generated ads are accurate, diverse, commercially valuable, and aligned with user interests.
  • Extensive offline and online experiments demonstrate the effectiveness of our approach. We observe substantial improvements in offline metrics such as HitRatio, as well as in online metrics like GMV.
@@ 2. Related Work
2.1. Sequential Recommendation
Sequential recommendation leverages the user behavior sequence on items (ads) to predict the next item that user is likely to click or convert on (Xu et al., 2019a; Zhou et al., 2022). Early approaches in sequential recommendation primarily focused on ID-based methods. These methods typically assign a unique ID to each item and employ sequential deep learning models, such as RNNs (Xu et al., 2019b), CNNs (Yan et al., 2019), and Transformers (Kang and McAuley, 2018; Sun et al., 2019), to learn sequence representations for next-item prediction. However, these approaches often struggle with issues related to cold start and data sparsity. To address these challenges, subsequent research has incorporated additional feature information, including categorical features, numerical features, and graph structure features (Kumar et al., 2019; Yuan et al., 2021; Chang et al., 2021; Liu et al., 2016). Some studies have also explored the integration of multi-modal information, such as item text descriptions and images (Li et al., 2023; Liu et al., 2021; Xie et al., 2022b). By leveraging this supplementary information, sequential recommendation methods can effectively mitigate cold start and data sparsity problems.
Moreover, some researchers have begun to adopt large language models (LLMs) to comprehend item features and infer user preferences, enabling direct recommendations without relying on traditional sequential recommendation methods (Bao et al., 2023; Harte et al., 2023; Wei et al., 2024; Bao et al., 2024). In this paper, we introduce LLMs into a large-scale display advertising system to tackle challenges encountered in industrial scenarios, such as implicit user intent and high computational costs.
@@ 2.2. Large Language Models-based Recommender
Large Language Models (LLMs) have demonstrated remarkable capabilities in both understanding and generation tasks (Yao et al., 2024; Huang et al., 2022), making them widely applicable across various domains, including document summarization (Koh et al., 2022; Zhang et al., 2024b), conversational agents (He et al., 2023; Wang et al., 2023), code completion (Guo et al., 2023; Zhang et al., 2024c), and others (Zhu et al., 2023; Zhai, 2024). Recently, researchers have begun exploring the application of LLMs in recommender systems (Wu et al., 2024). These efforts can be broadly categorized into two roles for LLMs: feature encoder and ranker.
As feature encoders, LLMs leverage their deep understanding of textual content to generate rich representations of users and items. Typically, user or item features are transformed into structured textual descriptions using pre-designed templates, which are then processed by LLMs to obtain feature embeddings (Ren et al., 2024; Wei et al., 2024; Bao et al., 2024). These embeddings can subsequently be utilized for downstream tasks such as candidate retrieval or ranking, improving personalization and recommendation quality.
As rankers, LLMs leverage their generative capabilities to predict user preferences based on historical behavior sequences. In this paradigm, user behavior sequences are represented as textual sequences, and LLMs are trained to predict the next likely item (or user action) based on context (Zhang et al., 2023; Harte et al., 2023). To adapt LLMs effectively for ranking tasks, fine-tuning on domain-specific recommendation datasets is typically required, ensuring the models acquire relevant domain knowledge and effectively capture user preferences (Bao et al., 2023; Harte et al., 2023; Lin et al., 2024).
Despite these advancements, the direct application of LLMs in industrial display advertising remains challenging due to several factors, including implicit user intent, high computational costs, and the need for real-time inference. In this work, we present the first application of LLM-based generative retrieval in an industrial display advertising system. To address these challenges, we introduce intent-aware prompt engineering and ad-specific knowledge alignment, enabling LLMs to generate diverse, business-aligned, and user-tailored ad recommendations. Furthermore, we mitigate computational overhead by deploying LLMs in a hybrid architecture that integrates latency-tolerant and latency-sensitive services. This approach ensures that the benefits of LLM-based retrieval can be realized in a scalable and cost-efficient manner.
@@ 2.3. Retrieval in Advertising Systems
Display advertising systems typically follow a multi-stage architecture composed of multiple stages, including retrieval, corse-ranking, ranking, re-ranking, and others(Cui et al., 2020; Gharibshah et al., 2020). The retrieval stage, positioned at the top of the funnel, plays a critical role in identifying user interests and generating a pool of candidate ads for further processing (Xie et al., 2022a; Ji et al., 2024). Most advertising systems adopt ID-based learning to ranking mechanisms to recall all ads that users may find appealing. These methods often rely on a two-tower model, where one tower encodes user features and the other encodes item (ad) features (Cen et al., 2020; Huang et al., 2020).
Display advertising systems typically follow a multi-stage architecture consisting of retrieval, coarse ranking, ranking, re-ranking, and others (Cui et al., 2020; Gharibshah et al., 2020). The retrieval stage, positioned at the entry of the process, is crucial for identifying user interests and generating an initial pool of candidate ads for subsequent ranking stages (Xie et al., 2022a; Ji et al., 2024). Most advertising systems employ ID-based learning-to-rank mechanisms to retrieve relevant ads for users. These methods often utilize a two-tower model, where one tower encodes user features and the other encodes item (ad) features (Cen et al., 2020; Huang et al., 2020). While efficient, ID-based retrieval approaches exhibit several limitations: (1) Information Cocoon Effect: The reliance on past interactions can reinforce existing preferences, leading to a lack of novelty in recommendations. (2) Limited Content Awareness: ID-based approaches often fail to leverage rich ad content, missing potential signals for better personalization. (3) Reduced Diversity: The retrieved candidate set may be dominated by frequently interacted items, limiting exposure to new and diverse ads.
To address these challenges, we propose incorporating LLMs into the retrieval process. Unlike traditional ID-based methods, LLMs can analyze contextual information and generate a more diverse set of candidate ads by expanding retrieval beyond strict collaborative filtering signals. By leveraging their generative capabilities and semantic understanding, LLMs enhance retrieval diversity, improve content-awareness, and ultimately increase user engagement and advertising effectiveness.
@@ 3.1. Problem Definition
Let $\mathcal{U}$ and $\mathcal{A}$ denote the sets of users and advertisements (ads), respectively, within the target advertising system. Given the limited user interactions with ads, we introduce a content domain to enrich user behavior data. Let $\mathcal{C}$ denote the set of content items in the system. The sequence of user behaviors is represented as $\mathcal{S}_{u}=[i^{1}_{u},i^{2}_{u},\dots,i^{L}_{u}]$ in chronological order, where $u\in\mathcal{U}$, $i_{u}\in\mathcal{A}\cup\mathcal{C}$, and $L=|\mathcal{S}_{u}|$ is the length of the user’s behavior sequence. Given this user behavior sequence, the goal of generative retrieval is to generate the next relevant ad, $a^{L+1}_{u}\in\mathcal{A}$, optimizing both user engagement and business objectives.
@@ 3.2. Ads Indexing
Traditional advertising systems typically index ads incrementally and learn dense embeddings for each ad (Wu et al., 2020; Chang et al., 2021). However, this indexing lacks semantic information, creating a gap between the advertising system and human understanding. To bridge this gap, we introduce Semantic IDs (S-IDs) based on ad features, inspired by previous works on item indexing (Rajput et al., 2024; Zheng et al., 2024; Geng et al., 2022).
To learn S-IDs, we first define a feature mapping rule in the form of a template string to generate text description of each ad, as follows:
@@
After obtaining the textual description $t_{a}$, we employ a pre-trained language model (e.g. Hunyuan (Sun et al., 2024), E5 (Wang et al., 2022)) to encode $t_{a}$, bridging the gap between the advertising system and human understandable text. This step can be formulated as
(1) $\mathbf{x}_{a}=f_{PLM}(t_{a};\theta_{PLM})$
where $f_{PLM}$ represents the pre-trained language model parameterized by $\theta_{PLM}$, and $\mathbf{x}_{a}\in\mathbb{R}^{d_{PLM}}$ is the textual embedding of ad $a$, and $d_{PLM}$ denotes the output dimension of the pre-trained language model.
After that, we employ a Residual-Quantized Variational AutoEncoder (RQ-VAE) (Lee et al., 2022) to generate concise semantic IDs for the ads. The RQ-VAE takes the ad’s embedding $\mathbf{x}_{a}$ as input and uses an encoder-quantization-decoder mechanism to generate a list of discrete semantic tokens. In the encoding phase, the textual embedding $\mathbf{x}_{a}$ is encoded by an encoder, $f_{En}(\cdot;\theta_{En})$, to obtain a hidden embedding, formulated as $\hat{\mathbf{z}}_{a}=f_{En}(\mathbf{x}_{a};\theta_{En})$, where $\hat{\mathbf{z}}_{a}\in\mathbb{R}^{d_{RQ}}$ denotes the hidden embedding, and $d_{RQ}$ is the hidden embedding’s dimension. During the quantization phase, we apply $M$ layers of residual vector quantization. For each layer $l$, we maintain a codebook $\mathcal{C}^{l}=\{\mathbf{e}_{k}^{l}\}_{k=1}^{K}$, where $\mathbf{e}_{k}^{l}\in\mathbb{R}^{d_{RQ}}$ and $K$ denotes the number of codes in the codebook. The quantization process is defined as follows:
(2) $c^{l}_{a}=$ $\text{argmin}_{k}||\mathbf{r}^{l}_{a}-\mathbf{e}_{k}^{l}||_{2}^{2}$
(3) $\mathbf{r}^{l+1}_{a}$ $=\mathbf{r}^{l}_{a}-\mathbf{e}_{c^{l}_{a}}^{l}$
(4) $\mathbf{z}_{a}$ $=\sum_{l=1}^{M}\mathbf{e}_{c^{l}_{a}}^{l}$
(5) $\mathcal{L}_{recons}=\sum_{a}||\mathbf{x}_{a}-\hat{\mathbf{x}}_{a}||_{2}^{2}$
(6) $\mathcal{L}_{quant}=\sum_{a}\sum_{l=1}^{M}\left(||\text{sg}[\mathbf{r}^{l}_{a}]-\mathbf{e}_{c^{l}_{a}}^{l}||_{2}^{2}+\beta_{quant}||\mathbf{r}^{l}_{a}-\text{sg}[\mathbf{e}_{c^{l}_{a}}^{l}]||_{2}^{2}\right)$
Algorithm 1 Trie-Tree Construction
1:Set of all ads’ S-IDs $\mathcal{C}_{A}=\{\mathbf{c}_{1},\mathbf{c}_{2},\dots,\mathbf{c}_{a},\dots,\mathbf{c}_{|\mathcal{A}|}\}$, where each S-IDs sequence $\mathbf{c}_{a}=[c_{a}^{1},c_{a}^{2},\dots,c_{a}^{l},\dots,c_{a}^{(M+1)}]$.
2:Trie-Tree $T$ representing all ads in $\mathcal{A}$.
3:Initialize: Create a head node $h$ for the trie-tree.
4:for each S-IDs sequence $\mathbf{c}_{a}\in\mathcal{C}_{A}$ do
5:    Set current node $cur\leftarrow h$
6:    for each S-ID $c_{a}^{l}$ in $\mathbf{c}_{a}$ do
7:         if $c_{a}^{l}$ is not a child of $c$ then
8:             Create a new node $n$
9:             Add $n$ as a child of $cur$ with label $c_{a}^{l}$
10:         end if
11:         Set $c\leftarrow$ child of $cur$ labeled by $c_{a}^{l}$
12:    end for
13:    Mark node $cur$ as the end of ad $a$
14:end for
15:return Trie $T$
This process ensures that ads are represented by discrete S-IDs, which allows the LLMs to generate valid ads from the predefined ads set during constrained decoding. In the ad indexing process, each ad $a$ is represented by a list of semantic tokens, denoted as $\hat{\mathbf{c}}_{a}=[c_{a}^{1},c_{a}^{2},\dots,c_{a}^{M}]$. To address potential collisions where different ads map to the same S-IDs, we introduce an additional code to distinguish them (Rajput et al., 2024). Consequently, the S-IDs for each ad are updated to $\mathbf{c}_{a}=[c_{a}^{1},c_{a}^{2},\dots,c_{a}^{L},c_{a}^{(M+1)}]$, where $c_{a}^{(M+1)}$ serves as the disambiguation code for ads with identical S-ID lists. To distinguish between different levels of semantic tokens, we use prefixes such as $<a\_,b\_,c\_,\cdots>$ for various S-ID levels.